# Reducing classical communication costs in multiplexed quantum repeaters using hardware-aware quasi-local policies

### v2

@@ IV.1 Comparison with doubling
Figure 6: Comparison of (top) average waiting times and (bottom) average age of the end-to-end link, for the FN swap-asap policy with and and without distillation. For the average waiting time, distill-swap outperforms the policy without distillation for low elementary link success probabilities, but it is not useful to distill when the success probability is high. At the same time, we see that it is never useful to swap-distill, both in terms of the average waiting time and the age of the end-to-end link. The CC costs of distillation are much higher when long virtual links are distilled, which provides intuition for the inefficacy of the swap-distill policy.
IV.2 Distillation-based policies
Now we present results for distillation-based policies when CC overheads are included. As mentioned earlier in Sec. II.1, the CC cost of distillation for any quasi-local policy is given by the length of longest links that are used for distillation. For local policies, distillation is only allowed at the elementary link level, and only when the links are freshly prepared. We continue using the distill-asap policy described in Sec. III. The CC overheads make the already probabilistic BBPSSW protocol even more costly. Furthermore, the CC costs of distilling longer (virtual) links is higher than distilling elementary links, and thus the question of whether to distill or not and whether to swap first or to distill first depends crucially on the CC overheads. We show the average waiting times for FN swap-asap policy without distillation and both distillation-based policies (swap-distill and distill-swap) in Fig. 6.
Figure 7: Comparison of (top) average waiting times and (bottom) average age of end-to-end link, distill-swap versions of FN swap-asap, SN swap-asap and the SN doubling policies when classical communication costs are included. FN swap-asap is the best policy for reducing the average waiting time and SN swap-asap for average age of the youngest end-to-end link.
Finally, in Fig. 7, we again compare the different distillation-based versions of quasi-local policies amongst themselves, viz., FN swap-asap, SN swap-asap and the SN doubling policies. We see that, with distillation (distill-swap version) and the relevant CC costs added, the trends remain similar to Fig. 5, i.e., FN swap-asap is still the best policy for reducing the average waiting time and SN swap-asap for average age of the youngest end-to-end link.
@@ IV.2 Distillation-based policies
V Unraveling the role of entanglement distillation policies for multiplexing
The BBPSSW entanglement distillation protocol considered in this work non-deterministically converts two links with low fidelity into one link with a higher fidelity. This immediately raises the following question: “Is distillation useful for all network parameter regimes?” Due to its non-deterministic nature, in certain parameter regimes, the slight increase in fidelity obtained by distillation might not be worth the potential loss of links, in case it fails. Furthermore, even without the consideration of CC costs, there exists an underlying asymmetry between the way entanglement swapping and distillation protocols work. In the former, the success probability is independent of the ages of the links, and thus it does not discriminate between longer (generally older) and shorter links, whereas the latter’s success crucially depends on the ages of the input links. This asymmetry leads to the the following question: “Should we distill and then swap or the other way around?” This consideration leads to two distinct policies, namely distill-swap and swap-distill, with different behaviours in different parameter regimes, as we show in Fig. 6. The aim of this section is to pursue the mechanisms behind these different behaviours and unravel them layer by layer. We make one simplifying assumption in this section, which is that classical communication costs associated with entanglement swapping and distillation are ignored. This is motivated by two reasons: firstly, for short chains, CC costs are usually ignored in existing literature on entanglement distribution policies [44, 86, 87, 88, 31], thus this assumption allows us to fairly compare the performance of our policies with those in the existing literature. Secondly, ignoring CC simplifies our model and makes some analyses more transparent. At the same time, the conclusions we come to in this section hold qualitatively even when CC is accounted for.
@@ V Unraveling the role of entanglement distillation policies for multiplexing
Another important consideration is that both links being distilled should be entangled. If one of the links is not entangled, then it is better to simply discard the unentangled link and keep the entangled one. We elaborate on this point in Appendix C. With this in mind, let us note that under the Pauli noise model that we consider, the fidelity according to Eq. (1) is bounded from below by $0.25$ (in the limit $m\to\infty$), and it is equal to approximately $0.5259$ when $m=m^{\star}$. The entanglement threshold for the states that we consider is $0.5$, meaning that the link is entangled if and only if its fidelity strictly exceeds $0.5$, and thus the age at which the link is no longer entangled is $m=\lceil m^{\star}\log(3)\rceil\approx\lceil 1.09m^{\star}\rceil$; see Appendix B for the details. This means that, depending on the value of $m^{\star}$, it can be useful to distill links that are older than $m^{\star}$ time steps. However, the probability of successful distillation falls with increasing ages of the links, and hence using a very old link in conjunction with a very young link increases the likelihood of wasting the high-quality young link. Thus, while distillation might be useful in this scenario, it is not necessarily worthwhile. Also, for the moderate values of $m^{\star}$ considered in our simulations, we choose to discard links at $m^{\star}$ time steps, instead of the strict limit of $\lceil m^{\star}\log(3)\rceil$, even when distillation is considered.
We now present simulation results for a seven-node chain with seven channels. (We also provide analytical results in Appendix D.) A somewhat large value of the cutoff is chosen ($m^{\star}=24$) to show some important scaling behaviour with increasing age $m_{0}$ of the fresh (newly created) links. Fig. 8 shows the average waiting time and average age of end-to-end links as a function of the ages of the elementary links $m_{0}$ when they are fresh, i.e., newly created. For both swap-distill and distill-swap, we have fixed the entanglement swapping policy to FN swap-asap.
Figure 8: Average waiting time (top) and average age of end-to-end links (bottom) as a function of the age $m_{0}$ of newly-created elementary links. The entanglement swapping policy is fixed to FN swap-asap. It is clear that swap-distill is the policy of choice in terms of reducing the waiting time when fresh elementary links have high fidelity (low age); otherwise, distill-swap performs better.
@@
Some estimates of the $m_{0}$-thresholds up to which the two distillation policies are effective can be made using some rather simple arguments. If the fresh elementary links have age $m_{0}$, the age of the distilled link saturates to a minimum value $m^{d}_{\min}(m_{0})$ which is a function of $m_{0}$ (see, e.g., the pumping policy in Appendix C). In order to get an end-to-end link via entanglement swapping of such distilled links, the following condition must be satisfied:
@@
Figure 9: Average waiting time improvement factor for distillation with the swap-distill and distill-swap policies (with FN swap-asap entanglement swapping). When $p_{\ell}$ is low, distill-swap performs better, and when $p_{\ell}$ is high swap-distill performs better. These are general trends and in concurrence with the Fig. 10 (center), where for low $p_{\ell}$ distill-swap is better but as $p_{\ell}$ increases swap-distill catches up. Furthermore, for large $p_{\ell}$, distillation is either not very useful, and can even be detrimental.
We further elaborate on our answer to the question of whether to distill first or to swap first in Fig. 9. To do this, we plot the waiting time improvement factor for distillation, which we define to be the ratio of the average waiting time without distillation to the average waiting time with distillation. Apart from the quality of fresh links (the parameter $m_{0}$), as shown in Fig. 8, the choice between the two distillation orderings also depends on the other network parameters. As an illustration, we show that for the parameters $(n,n_{ch},m^{\star},p_{\ell})$ chosen in Fig. 9, when $p_{\ell}$ is low, distill-swap performs better, and when $p_{\ell}$ is high swap-distill performs better. This observation can be understood in light of the analytical results in Appendix D and Fig. 8. Indeed, we observe that for low $p_{\ell}$, links are typically older when they are ready to be swapped, and we have seen in the previous discussions in this section that distill-swap performs better in this case, because long entanglement swaps can only be possible once the ages of the participating links have been reasonably reduced. Also, distilling after entanglement swapping is highly improbable to be successful because the ages are going to only increase due to the swapping. On the other hand, when links have a high $p_{\ell}$, swap-distill performs better, because we observe that such links are typically younger when they are ready to be swapped, and they are much more conducive to successful swapping. Furthermore, in this case, the non-deterministic swaps become the rate determining step, and hence having more opportunities to perform entanglement swapping is more important. Distilling links first reduces the number of such opportunities. Furthermore, we see that for large $p_{\ell}$, distillation is either not very useful or even detrimental. Note that these trends were also qualitatively seen in Fig. 6, when CC costs were accounted for.
@@
Figure 10: Average waiting time improvement factor for distillation with the swap-distill and distill-swap policies (with FN swap-asap entanglement swapping). Distillation is not very useful, in fact mostly detrimental, when (top) fresh elementary links are perfect ($m_{0}=0$, $100\%$ fidelity), but when fresh elementary links are imperfect (center) ($m_{0}=1$, approx. $90\%$ fidelity), a dramatic reduction in waiting time can be obtained by performing distillation. (Bottom) Improvement factor for average age of youngest end-to-end link using distillation over non-distillation based (FN) policy. swap-distill provides negligible advantage but distill-swap provides substantial advantage.
@@
The discussion above leads us to the next question pertaining to the usefulness of distillation-based policies. We are yet to show whether there is any advantage in performing distillation. Distillation is a probabilistic process, and even when it succeeds it leads to a reduction in the number of active links. Hence, it is not obvious that a policy without distillation, such SN or FN swap-asap alone, cannot perform better than its distillation-based counterpart. Let us look at two cases: when fresh elementary links are perfect, i.e., $m_{0}=0$, and when they are imperfect, say $m_{0}=1$ for concreteness. In Fig. 10 we look at a seven-node chain with $n_{ch}=12$, $p_{sw}=0.5$, and $m^{\star}=8$ ($m_{0}=1$ corresponds to roughly $90\%$ fidelity in this case). We choose FN swap-asap as our swapping policy. We show the average waiting time improvement factor for distillation-based policies, defined as the ratio of average waiting time without distillation and with distillation as a function of the elementary link success probability $p_{\ell}$. An improvement factor of greater than 1 indicates an advantage in using distillation and vice versa. We see in Fig. 10 (top) that distillation is mostly detrimental when freshly produced elementary links are perfect (100% fidelity). Both swap-distill and distill-swap increase the average waiting time, for all values of the elementary link probability. Further, the disadvantage is higher in the case of distill-swap. This can understood by the following observation. The primary role of distillation in the case when elementary links are perfect is to reduce the age of long or older links obtained as a consequence of waiting for other links and/or via entanglement swaps. The longest links tend to be the oldest ones in the network, due to the $m_{1}+m_{2}$ age-update rule for entanglement swapping. If we distill first, then initially, since most links are perfect, distillation is never invoked and after entanglement swaps, when distillation is needed, we have to wait an extra time step before distillation is attempted. This extra time step reduces the success probability of distillation of already old links, following Eq. (2). This effect is even more magnified when CC costs are added, since swaps will now lead to even older links. This is also the reason why swap-distill has some positive improvement in some parameter regimes. Although not shown in Fig. 10, there exists a small parameter regime of low $m^{\star},p_{\ell}$ and $n_{ch}$ in which swap-distill leads to a small positive improvement (3-5%). The choice of parameters in the figure is motivated by the intent to show the most prominent and overarching trends. On the other hand, when elementary links are imperfect, distillation is essential in increasing the scale, fidelity and throughput of the network, indicated by the significant reduction in waiting time seen in Fig. 10 (bottom). When the elementary link probability is low, distill-swap performs much better, and when it is is high, swap-distill catches up in terms of improvement. This is in agreement with the results shown in Fig. 8 and 9. Although, the trends of improvement factor variation with increasing elementary link probability change between Fig. 9 and 10. In terms of the end-to-end link fidelity swap-distill provides negligible advantage but distill-swap provides substantial advantage as shown in Fig. 10 (bottom).
V.1 Main messages for distillation
@@
    Yes (generally), when fresh elementary links are not perfect. In this case, distillation leads to a significant improvement in both the average waiting time and the average age of the end-to-end link. Although this choice is not straightforward, in extremely resource constrained scenarios, such as very low $p_{\ell}$ or $p_{sw}$ it might be disadvantageous to distill (see Sec. VI.2), even when fresh links are imperfect.
@@
    If links are have low $p_{\ell}$ or low initial fidelity, distilling first is more advantageous, in the opposite case it is better to swap first.
@@
VI Experimental Implementation
Figure 11: (a) Proposed implementation of a multiplexed repeater chain using a high-dimensional biphoton frequency comb (BFC). (b) Measured frequency correlation matrix of the BFC for the multiplexed elementary link that we implemented. (c) Polarization entanglement measurements of the implemented BFC. (d) Schmidt number and Hilbert-space dimensionality scaling for different dimensions in the implemented BFC.
Based on the multiplexed model introduced in Fig. 1, in this section we propose a proof-of-principle experimental realization of a quantum network with multiplexing using a high-dimensional biphoton frequency comb (BFC) in Fig. 11.
As part of this proposal, we present results of an implementation a multiplexed elementary link using a high-dimensional BFC. First, the BFC state is prepared by passing the spontaneous parametric down-conversion (SPDC) photons generated from a type-II periodically-poled KTiOPO4 (ppKTP) waveguide through a fiber Fabry–Pérot cavity (FFPC) [89, 90]. Due to the type-II phase-matching condition, the signal and idler photons can be efficiently separated by a polarization beam-splitter (PBS), enabling deterministic BFC generation without post-selection. The signal and idler photons are then temporally overlapped at a fiber beamsplitter with orthogonal polarizations, by using a tunable delay line (DL) and polarization controllers [89]. This configuration generates polarization entangled BFC state
@@ VI Experimental Implementation
We perform the Schmidt mode decomposition of the measured joint spectral intensity [89], and summarize the extracted Schmidt mode number ($K$), and the corresponding Hilbert space dimension ($K^{2}$). When reconstructing the density matrix, we observe that the Schmidt mode number and Hilbert-space dimension scales with the dimension $d$ of the biphoton frequency comb, as shown in Fig. 11(d). The Schmidt mode number is $2.98$ with $d=3$, and $K=7.03$ when we increase the dimension to $d=9$. The deviation from ideal scaling in Fig. 11(d) mainly comes from the sinc function of the SPDC spectrum, which can be circumvented by using a flat-top broadband SPDC source. The Hilbert-space dimension can be further increased by using a SPDC source with broader spectrum bandwidth and FFPC with smaller FSR. We also note that by passing only the signal photons through a FFPC, the idler photons will still exhibit a comb-like behavior [91, 92]. Such a scheme can provide higher photon flux due to less filtering loss, and with potentially higher secure key rates for quantum key distribution [93, 94, 95, 96]. Although we have already experimentally implemented the multiplexed source required for the proposed experimental implementation, integration of the scheme with quantum memories, implementing entanglement swapping, and distillation of links has not been done yet. At the same time, there has been great progress in each of these key steps for the construction of a long-distance quantum network.
VI.1 Calculation of elementary link probability, time step, and memory cutoff
In this section, we show how to translate physical parameters into the values of $p_{\ell}$, $m^{\star}$, and $\Delta t$ within our theoretical framework
@@ VI.1 Calculation of elementary link probability, time step, and memory cutoff
where $\eta_{\ell}=\exp(-\frac{\ell}{12\text{ km}})$ is the photon-loss contribution, with $\ell$ being the length of the elementary link. For =40 $\ell=40\text{ km}$, we have $\eta_{\ell}\approx 0.036$. The factor of 112 $\frac{1}{12\text{ km}}$ in the exponent corresponds to the 3.63 dB loss in the 10 km channel reported in the experiment in Ref. [92]. In addition, $\eta=0.69$ is the Debye–Waller factor, the ideal efficiency of the memory (fraction of photons captured in the memory out of all the incident photons) [97], for a rare-Earth metal based quantum memory [98], and $\eta_{r}=0.79$ (approximately 1 dB) is the loss from other residual factors, which could include detector efficiencies, optical component loss, etc. [89]. Altogether, we have that $p_{\ell}=\eta_{r}\eta_{\ell}\eta^{2}\approx 0.0134$.
Heralding time step $\Delta t$.
This is given by the elementary link distance $\ell$ as
$\Delta t=\frac{n\ell}{c}\,\text{seconds},$ (10)
where $n$ is the refractive index of the communication medium, $\ell$ is length of the elementary link, and $c$ is the speed of light in vacuum.
@@
In Table 1, we calculate our model parameters for various values of the elementary link distance.
$\ell$ (km) $p_{\ell}$ $\Delta t$ (ms) $m^{\star}$
25.0 0.05 0.125 8
15.0 0.11 0.075 13
5.0 0.25 0.025 40
Table 1: Parameters $p_{\ell}$, $\Delta t$, and $m^{\star}$ of our theoretical model, calculated for varying elementary link distances $\ell$ based on the physical aspects of the quantum channels and memories.
@@
VI.2 Figures of merit and repeater chain design for existing memory platforms

### v3

@@ Introduction
Figure 6: Performance of various quasi-local policies with distillation. Comparison of average waiting times (a) and average age of the end-to-end link (b) for the FN swap-asap policy with and without distillation. For the average waiting time, distill-swap outperforms the policy without distillation for low elementary link success probabilities, but it is not useful to distill when the success probability is high. At the same time, we see that it is never useful to swap-distill, both in terms of the average waiting time and the age of the end-to-end link. The CC costs of distillation are much higher when long virtual links are distilled, which provides intuition for the inefficacy of the swap-distill policy. Average waiting times (c) and average age of end-to-end link (d) for the distill-swap versions of FN swap-asap, SN swap-asap and the SN doubling policies when classical communication costs are included. FN swap-asap is the best policy for reducing the average waiting time and SN swap-asap for average age of the youngest end-to-end link.
Distillation-based policies
Now we present results for distillation-based policies when CC overheads are included. As mentioned earlier in Classical communication, the CC cost of distillation for any quasi-local policy is given by the length of longest links that are used for distillation. For local policies, distillation is only allowed at the elementary link level, and only when the links are freshly prepared. We continue using the distill-asap policy described in Quasi-local multiplexing policies. The CC overheads make the already probabilistic BBPSSW protocol even more costly. Furthermore, the CC costs of distilling longer (virtual) links is higher than distilling elementary links, and thus the question of whether to distill or not and whether to swap first or to distill first depends crucially on the CC overheads. We show the average waiting times for FN swap-asap policy without distillation and both distillation-based policies (swap-distill and distill-swap) in Fig. 6(a,b).
Finally, in Fig. 6(c,d), we again compare the different distillation-based versions of quasi-local policies amongst themselves, viz., FN swap-asap, SN swap-asap and the SN doubling policies. We see that, with distillation (distill-swap version) and the relevant CC costs added, the trends remain similar to Fig. 5, i.e., FN swap-asap is still the best policy for reducing the average waiting time and SN swap-asap for average age of the youngest end-to-end link.
@@
Unraveling the role of entanglement distillation policies for multiplexing
The BBPSSW entanglement distillation protocol considered in this work non-deterministically converts two links with low fidelity into one link with a higher fidelity. This immediately raises the following question: “Is distillation useful for all network parameter regimes?” Due to its non-deterministic nature, in certain parameter regimes, the slight increase in fidelity obtained by distillation might not be worth the potential loss of links, in case it fails. Furthermore, even without the consideration of CC costs, there exists an underlying asymmetry between the way entanglement swapping and distillation protocols work. In the former, the success probability is independent of the ages of the links, and thus it does not discriminate between longer (generally older) and shorter links, whereas the latter’s success crucially depends on the ages of the input links. This asymmetry leads to the the following question: “Should we distill and then swap or the other way around?” This consideration leads to two distinct policies, namely distill-swap and swap-distill, with different behaviors in different parameter regimes, as we show in Fig. 6. The aim of this section is to pursue the mechanisms behind these different behaviors and unravel them layer by layer. We make one simplifying assumption in this section, which is that classical communication costs associated with entanglement swapping and distillation are ignored. This is motivated by two reasons: firstly, for short chains, CC costs are usually ignored in existing literature on entanglement distribution policies [44, RL22, 90, 91, 31], thus this assumption allows us to fairly compare the performance of our policies with those in the existing literature. Secondly, ignoring CC simplifies our model and makes some analyses more transparent. At the same time, the conclusions we come to in this section hold qualitatively even when CC is accounted for.
@@
Another important consideration is that both links being distilled should be entangled. If one of the links is not entangled, then it is better to simply discard the unentangled link and keep the entangled one. We elaborate on this point in “Methods” when discussing the BBPSSW distillation protocol. With this in mind, let us note that under the Pauli noise model that we consider, the fidelity according to Eq. (1) is bounded from below by $0.25$ (in the limit $m\to\infty$), and it is equal to approximately $0.5259$ when $m=m^{\star}$. The entanglement threshold for the states that we consider is $0.5$, meaning that the link is entangled if and only if its fidelity strictly exceeds $0.5$, and thus the age at which the link is no longer entangled is $m=\lceil m^{\star}\log(3)\rceil\approx\lceil 1.09m^{\star}\rceil$ (also see “Methods”). This means that, depending on the value of $m^{\star}$, it can be useful to distill links that are older than $m^{\star}$ time steps. However, the probability of successful distillation falls with increasing ages of the links, and hence using a very old link in conjunction with a very young link increases the likelihood of wasting the high-quality young link. Thus, while distillation might be useful in this scenario, it is not necessarily worthwhile. Also, for the moderate values of $m^{\star}$ considered in our simulations, we choose to discard links at $m^{\star}$ time steps, instead of the strict limit of $\lceil m^{\star}\log(3)\rceil$, even when distillation is considered.
We now present simulation results for a seven-node chain with seven channels. We also provide analytical results for some simpler cases in Supplementary Note 3. A somewhat large value of the cutoff is chosen ($m^{\star}=24$) to show some important scaling behaviour with increasing age $m_{0}$ of the fresh (newly created) links. Fig. 7 shows the average waiting time and average age of end-to-end links as a function of the ages of the elementary links $m_{0}$ when they are fresh, i.e., newly created. For both swap-distill and distill-swap, we have fixed the entanglement swapping policy to FN swap-asap.
Figure 7: Should we swap-then-distill or distill-then-swap? Average waiting time (a) and average age of end-to-end links (b) as a function of the age $m_{0}$ of newly-created elementary links. The entanglement swapping policy is fixed to FN swap-asap. It is clear that swap-distill is the policy of choice in terms of reducing the waiting time when fresh elementary links have high fidelity (low age); otherwise, distill-swap performs better.
@@
Some estimates of the $m_{0}$-thresholds up to which the two distillation policies are effective can be made using some rather simple arguments. If the fresh elementary links have age $m_{0}$, the age of the distilled link saturates to a minimum value $m^{d}_{\min}(m_{0})$ which is a function of $m_{0}$ (see, e.g., the pumping policy in Supplementary Note 2). In order to get an end-to-end link via entanglement swapping of such distilled links, the following condition must be satisfied:
@@
Figure 8: Is distillation useful? Average waiting time improvement factor for distillation with the swap-distill and distill-swap policies (with FN swap-asap entanglement swapping). (a) When $p_{\ell}$ is low, distill-swap performs better, and when $p_{\ell}$ is high swap-distill performs better. These are general trends and in concurrence with (c), where for low $p_{\ell}$ distill-swap is better but as $p_{\ell}$ increases swap-distill catches up. Furthermore, for large $p_{\ell}$, distillation is either not very useful, and can even be detrimental. Distillation is not very useful, in fact mostly detrimental, when (b) fresh elementary links are perfect ($m_{0}=0$, $100\%$ fidelity), but when fresh elementary links are imperfect (c) ($m_{0}=1$, approx. $90\%$ fidelity), a dramatic reduction in waiting time can be obtained by performing distillation. (d) Improvement factor for average age of youngest end-to-end link using distillation over non-distillation based (FN) policy. swap-distill provides negligible advantage but distill-swap provides substantial advantage.
We further elaborate on our answer to the question of whether to distill first or to swap first, and whether or not distillation is useful in Fig. 8. To do this, we plot the waiting time improvement factor for distillation, which we define to be the ratio of the average waiting time without distillation to the average waiting time with distillation. Apart from the quality of fresh links (the parameter $m_{0}$), as shown in Fig. 7, the choice between the two distillation orderings also depends on the other network parameters. As an illustration, we show that for the parameters $(n,n_{ch},m^{\star},p_{\ell})$ chosen in Fig. 8 (a), when $p_{\ell}$ is low, distill-swap performs better, and when $p_{\ell}$ is high swap-distill performs better. This observation can be understood in light of the analytical results in Supplementary Note 3 and Fig. 7. Indeed, we observe that for low $p_{\ell}$, links are typically older when they are ready to be swapped, and we have seen in the previous discussions in this section that distill-swap performs better in this case, because long entanglement swaps can only be possible once the ages of the participating links have been reasonably reduced. Also, distilling after entanglement swapping is highly improbable to be successful because the ages are going to only increase due to the swapping. On the other hand, when links have a high $p_{\ell}$, swap-distill performs better, because we observe that such links are typically younger when they are ready to be swapped, and they are much more conducive to successful swapping. Furthermore, in this case, the non-deterministic swaps become the rate determining step, and hence having more opportunities to perform entanglement swapping is more important. Distilling links first reduces the number of such opportunities. Furthermore, we see that for large $p_{\ell}$, distillation is either not very useful or even detrimental. Note that these trends were also qualitatively seen in Fig. 6, when CC costs were accounted for.
@@
The discussion above leads us to further explore the question pertaining to the usefulness of distillation-based policies. Distillation is a probabilistic process, and even when it succeeds it leads to a reduction in the number of active links. Hence, it is not obvious that a policy without distillation, such SN or FN swap-asap alone, cannot perform better than its distillation-based counterpart. Let us look at two cases: when fresh elementary links are perfect, i.e., $m_{0}=0$, and when they are imperfect, say $m_{0}=1$ for concreteness. In Fig. 8(b–d) we look at a seven-node chain with $n_{ch}=12$, $p_{sw}=0.5$, and $m^{\star}=8$ ($m_{0}=1$ corresponds to roughly $90\%$ fidelity in this case). We choose FN swap-asap as our swapping policy. We show the average waiting time improvement factor for distillation-based policies, defined as the ratio of average waiting time without distillation and with distillation as a function of the elementary link success probability $p_{\ell}$. An improvement factor of greater than 1 indicates an advantage in using distillation and vice versa. We see in Fig. 8(b) that distillation is mostly detrimental when freshly produced elementary links are perfect (100% fidelity). Both swap-distill and distill-swap increase the average waiting time, for all values of the elementary link probability. Further, the disadvantage is higher in the case of distill-swap. This can understood by the following observation. The primary role of distillation in the case when elementary links are perfect is to reduce the age of long or older links obtained as a consequence of waiting for other links and/or via entanglement swaps. The longest links tend to be the oldest ones in the network, due to the $m_{1}+m_{2}$ age-update rule for entanglement swapping. If we distill first, then initially, since most links are perfect, distillation is never invoked and after entanglement swaps, when distillation is needed, we have to wait an extra time step before distillation is attempted. This extra time step reduces the success probability of distillation of already old links, following Eq. (35). This effect is even more magnified when CC costs are added, since swaps will now lead to even older links. This is also the reason why swap-distill has some positive improvement in some parameter regimes. Although not shown in Fig. 8(b), there exists a small parameter regime of low $m^{\star},p_{\ell}$ and $n_{ch}$ in which swap-distill leads to a small positive improvement (3-5%). The choice of parameters in the figure is motivated by the intent to show the most prominent and overarching trends. On the other hand, when elementary links are imperfect, distillation is essential in increasing the scale, fidelity and throughput of the network, indicated by the significant reduction in waiting time seen in Fig. 8(c). When the elementary link probability is low, distill-swap performs much better, and when it is is high, swap-distill catches up in terms of improvement. This is in agreement with the results shown in Fig. 7 and 8(a). In terms of the end-to-end link fidelity swap-distill provides negligible advantage but distill-swap provides substantial advantage as shown in Fig. 8(d). It is noteworthy that the trends of improvement factor variation with increasing elementary link probability change between Fig. 8(a) and Fig. 8(c) substantially with a small change in $m^{\star}$ from $8$ to $9$. This happens because $m^{\star}=9$ lies at the threshold near which distillation becomes beneficial, hence the high sensitivity to $m^{\star}$. As an aid to intuition, recall that if no distillation is used, then to establish an end-to-end link, when newly-created links have age $m_{0}=1$ in a chain of seven nodes, the smallest cutoff time that ensures an end-to-end entangled link is $m^{\star}=6$, and note that $m^{\star}=8,9$ are very close to this limit.
Main messages for distillation
@@
    Yes (generally), when fresh elementary links are not perfect. In this case, distillation leads to a significant improvement in both the average waiting time and the average age of the end-to-end link. Nonetheless, the decision to distill links is not straightforward, since in extremely resource constrained scenarios, such as very low $p_{\ell}$ or $p_{sw}$ it might again become disadvantageous to distill (see Figures of merit and repeater chain design for existing memory platforms), even when fresh links are imperfect.
@@
    If links have low $p_{\ell}$ or low initial fidelity, distilling first is more advantageous, in the opposite case it is better to swap first.
@@
Proposed experimental implementation
Figure 9: Proposed experimental implementation of a multiplexed repeater chain. (a) Proposed implementation of a multiplexed repeater chain using a high-dimensional biphoton frequency comb (BFC). (b) Measured frequency correlation matrix of the BFC for the multiplexed elementary link that we implemented. (c) Polarization entanglement measurements of the implemented BFC. (d) Schmidt number and Hilbert-space dimensionality scaling for different dimensions in the implemented BFC.
Based on the multiplexed model introduced in Fig. 1, in this section we propose a proof-of-principle experimental proposal of a quantum network with multiplexing using a high-dimensional biphoton frequency comb (BFC) in Fig. 9.
As part of this proposal, we present results of an implementation of a multiplexed elementary link using a high-dimensional BFC. First, the BFC state is prepared by passing the spontaneous parametric down-conversion (SPDC) photons generated from a type-II periodically-poled KTiOPO4 (ppKTP) waveguide through a fiber Fabry–Pérot cavity (FFPC) [92, 93, 94]. Due to the type-II phase-matching condition, the signal and idler photons can be efficiently separated by a polarization beam-splitter (PBS), enabling deterministic BFC generation without post-selection. The signal and idler photons are then temporally overlapped at a fiber beamsplitter with orthogonal polarizations, by using a tunable delay line (DL) and polarization controllers [92]. This configuration generates polarization entangled BFC state
@@
We perform the Schmidt mode decomposition of the measured joint spectral intensity [92], and summarize the extracted Schmidt mode number ($K$), and the corresponding Hilbert space dimension ($K^{2}$). In particular, $K$ represents the effective orthogonal modes in the system. When reconstructing the density matrix, we lower bound the Schmidt mode number, which scales linearly with the dimension $d$ of the biphoton frequency comb. The Hilbert-space dimensionality can then be estimated by $K^{2}$. In Fig. 9(d), we plot the Schmidt mode number and the Hilbert-space dimension as a function of the dimension $d$, and we compare it with the ideal, theoretical scaling. We observe that the Schmidt mode number is $K=2.98$ with $d=3$, and $K=7.03$ when we increase the dimension to 9. Such BFC state provides up to seven effective frequency modes for multiplexing. The deviation from ideal scaling in Fig. 9(d) mainly comes from the falloff from the sinc function of the SPDC spectrum, which can be circumvented by using a flat-top broadband SPDC source. The Hilbert-space dimension can be further increased by using a SPDC source with broader spectrum bandwidth and FFPC with smaller FSR. We also note that by passing only the signal photons through a FFPC, the idler photons will still exhibit a comb-like behavior [96, 95]. Such a scheme can provide higher photon flux due to less filtering loss, and with potentially higher secure key rates for quantum key distribution [97, 98, 99, 100]. Although we have already experimentally implemented the multiplexed source required for the proposed experimental implementation, integration of the scheme with quantum memories, implementing entanglement swapping, and distillation of links has not been done yet. At the same time, there has been great progress in each of these key steps for the construction of a long-distance quantum network.
Calculation of model parameters
In this section, we show how to translate physical parameters into the values of $p_{\ell}$, $m^{\star}$, and $\Delta t$ within our theoretical framework.
@@
where $\eta_{\ell}=\exp(-\frac{\ell}{12\text{ km}})$ is the photon-loss contribution, with $\ell$ being the length of the elementary link. For =40 $\ell=40\text{ km}$, we have $\eta_{\ell}\approx 0.036$. The factor of 112 $\frac{1}{12\text{ km}}$ in the exponent corresponds to the 3.63 dB loss in the 10 km channel reported in the experiment in Ref. [101]. In addition, $\eta=0.69$ is the Debye–Waller factor, the ideal efficiency of the memory (fraction of photons captured in the memory out of all the incident photons) [102], for a rare-Earth metal based quantum memory [103], and $\eta_{r}=0.79$ (approximately 1 dB) is the loss from other residual factors, which could include detector efficiencies, optical component loss, etc. [92]. Altogether, we have that $p_{\ell}=\eta_{r}\eta_{\ell}\eta^{2}\approx 0.0134$. In the case of emissive memories like those based on Diamond vacancies, an effective Debye–Waller factor of 0.5 is assumed, since a linear-optical Bell state measurement is needed at each node to teleport the state of the incoming SPDC photon into the memory.
Simulation time step $\Delta t$.
This is given by the elementary link distance $\ell$ and the source rate per channel as
$\Delta t=\max\Big{\{}\frac{1}{R},\frac{n\ell}{c}\Big{\}}\,\text{seconds},$ (10)
where $n$ is the refractive index of the communication medium, $\ell$ is length of the elementary link, and $c$ is the speed of light in vacuum and $R$ is the rate of ebit pairs generated by the dimmest channel of the multiplexed SPDC source (see Fig. 9) (b).
@@
Fresh elementary link age $m_{0}$.
This is given by the fidelity $f_{s}$ of the polarization entangled state produced by the SPDC source (which we assume to be the same for all channels, for simplicity) and the fidelity $f_{0}$ of the entangled state when it is freshly absorbed (emitted) by the quantum memory. The fidelity of the elementary link is thus given by $f_{e}\coloneqq f_{s}f_{0}^{2}$. This can be converted to the initial age of the elementary links using Eq. (26). Note that this also depends on $m^{\star}$.
For example, if the elementary link distance $\ell$ is varied from $5$ km to $25$ km, $p_{\ell}$ varies from $0.25$ to $0.05$, and $\Delta t$ goes from $0.025$ ms to $0.125$ ms. The latter in turn implies that $m^{\star}$ varies from $40$ to $8$, assuming a $T_{2}$ of approximately $1$ ms.
@@
Figures of merit and repeater chain design for existing memory platforms